two passenger trains A and b ,300km apart start towards each other at the same time.they meet after 2 hours .if train B travels 8/7 as fast as train A.What is the speed of each train
Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

They meet in two hours, and their combined speed is 300/2 or 150 kph.
Train A has speed x. Train B has speed (8x/7)
(8x/7)+(7x/7)=150
15x/7=150
15x=1050
x=70 kph Train A
8/7 x=80 kph Train B

Answer by stanbon(75887)

two passenger trains A and b ,300km apart start towards each other at the same time.they meet after 2 hours .if train B travels 8/7 as fast as train A.What is the speed of each train
-----
slower train DATA::
rate = x km/h ; time = 2 hrs ; distance = 2x km
--------------------
faster train DATA:
rate = (8/7)x km/h ; time = 2 hrs ; distance = 2(8/7)x = (16/7)x km
=======
Equation:
dist + dist = 300 km
2x + (16/7)x = 300
----
(30/7)x = 300
x = 70 km/h (slower train speed)
(8/7)x = 80 km/h (faster train speed)
==============
Cheers,
Stan H.
------------